# Ergodic Estimates of One-Step Numerical Approximations for Superlinear SODEs

### v1

@@ Abstract.
This paper establishes the first-order convergence rate for the ergodic error of numerical approximations to a class of stochastic ODEs (SODEs) with superlinear coefficients and multiplicative noise. By leveraging the generator approach to the Stein method, we derive a general error representation formula for one-step numerical schemes. Under suitable dissipativity and smoothness conditions, we prove that the error between the accurate invariant measure $\pi$ and the numerical invariant measure $\pi_{\tau}$ is of order $\mathscr{O}(\tau)$, which is sharp. Our framework applies to several recently studied schemes, including the tamed Euler, projected Euler, and backward Euler methods.
@@
Ergodic estimate, Stein method, stochastic one-step approximations, superlinear SODEs
@@ 1. Introduction
Here $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are measurable functions satisfying certain monotone and polynomial growth conditions.
@@
Under appropriate conditions, Eq. (1.1) admits a unique invariant measure $\pi$. While characterizing $\pi$ explicitly is generally infeasible, it is known only for specific cases, such as the gradient Langevin equation with additive noise or one-dimensional SODEs with multiplicative nondegenerate noise [5]. Therefore, constructing accurate numerical approximations of $\pi$ is a problem of significant interest. The study of long-time approximations for SODEs is a growing interest, with significant applications in scientific computing, statistics, machine learning, and generative AI. For instance, sampling in score-based diffusion models–a key technique in generative AI–fundamentally relies on long-time SODE approximations.
@@
The Euler–Maruyama (EM) scheme is effective for Lipschitz Eq. (1.1) (see, e.g., [2, 16]) or Markovian switching diffusion (see, e.g., [17]). However, for coefficients with superlinear growth, the EM scheme would blow up in $p$-th moment for all $p\geq 2$ [3], precluding ergodicity. Consequently, several modified schemes have been developed to preserve ergodicity, including the backward Euler method (BEM) and stochastic theta method [7, 10, 14], and the modified Euler method (MEM)–including tamed Euler method (TEM) and projected Euler method (PEM) [12, 13]; see also [6] for Markovian switching diffusions.
@@
A central question is to quantify the ergodic error $|\pi(\varphi)-\pi_{\tau}(\varphi)|$ between $\pi_{\tau}$ and $\pi$ for a given step-size $\tau\in(0,1)$ and a test function $\varphi$. Using the generator approach [1] to the Stein method [18], the authors of [2] derived an ergodic error representation for additive noise cases and obtained a rate of $\mathscr{O}(\tau^{1/2}|\ln\tau|)$ for the EM scheme (the so-called unadjusted Langevin algorithm in the Langenvin equation case) with Lipschitz drift and nondegenerate additive noise, where $\varphi$ is a Lipschitz function. This was later generalized in [4] to dissipative Lipschitz systems with nondegenerate multiplicative noise and jump diffusions, removing the logarithmic factor for test functions in $\mathcal{C}_{b}^{3}(\mathbb{R}^{d})$. We also refer to [15] and references therein for nonasymptotic bound analysis for the decreasing-step EM scheme.
@@
Since the weak convergence rate is typically twice the strong convergence rate, an ergodic error of order $\mathscr{O}(\tau)$ is expected. This paper achieves this optimal first-order rate for a family of superlinear SODEs with multiplicative noise and a class of one-step numerical approximations, assuming test functions in $\mathcal{C}_{b}^{4}(\mathbb{R}^{d})$.
Our contributions:
  1. (1)
    We derive the Stein equation with an explicit solution linked to the invariant measure under fewer conditions and introduce a new representation of the ergodic error for general one-step numerical approximations, decomposing it into an initial error and a remainder from the continuous interpolation (see Theorems 2.1 and 2.3, respectively).
  2. (2)
    Our framework applies to several recently studied numerical schemes–including the tamed Euler, projected Euler, and backward Euler methods–and establishes the optimal first order of convergence (see Theorem 3.5).
@@
The paper is organized as follows. Section 2 presents the main ideas, including the Stein equation and the general ergodic error estimate. Section 3 applies these results to superlinear SODEs and provides detailed analyses for specific numerical schemes.
@@ 2. Main Idea and Results
This section will introduce some frequently used notations, the Stein equation with its solution, and the main result on the ergodic error for general one-step approximations.
@@ 2.1. Notations
Let us begin with some frequently used notations. Throughout this paper, we denote by $\left\lvert\cdot\right\rvert$, $\langle\cdot,\cdot\rangle$, and $\left\lVert\cdot\right\rVert_{HS}$ the Euclidean norm, the inner product in $\mathbb{R}^{d}$, and the Hilbert–Schmidt norm in $\mathbb{R}^{d\times m}$.
In addition, we use $\mathcal{B}_{b}(\mathbb{R}^{d})$ and ${\rm Lip}(\mathbb{R}^{d})$ to denote the Banach spaces of all bounded measurable and Lipschitz continuous mappings $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, for $k\in\mathbb{N}_{+}$, denote by $C_{b}^{k}(\mathbb{R}^{d})$ the Banach spaces of all continuous functions with bounded partial derivatives $\nabla^{i}\varphi(x)\in\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R})$, the $i$-fold linear functional in $\mathbb{R}^{d}$ with $1\leq i\leq k$, endowed with the seminorm
@@ 2.2. Stein Equation
Assume $(X_{t}^{\cdot})_{t\geq 0}$ has bounded algebraic moments in the sense
@@
and $\mathbb{E}\int_{0}^{T}|\sigma_{j}(X_{t}^{\cdot})|^{2}\mathrm{d}t<\infty$ for all $T>0$ and $j=\{1,2,\cdots,m\}$, so that $\int_{0}^{\cdot}\sigma(X_{s}^{\cdot})\mathrm{d}W(s)$ is a martingale. We note that the definition (2.1) makes sense for $\varphi\in{\rm Lip}(\mathbb{R}^{d})$. Assume $(P_{t})_{t\geq 0}$ has a unique invariant measure $\pi$ with bounded algebraic moments and mixing property:
@@
For an $f\in C^{2}(\mathbb{R}^{d})$, Itô formula gives
@@
If $X_{0}\sim\pi$, we have $\int_{0}^{t}\mathbb{E}[\mathcal{A}f(X_{s})]\mathrm{d}s=0$ for any $t\geq 0$. Thus $\mathbb{E}[\mathcal{A}f(X_{0})]=0$ by the continuity of the process $X$.
@@
so that $f_{\varphi}$ given in (2.7) is a solution of the Stein equation (2.6). ∎
@@
The solution of the Stein equation (2.6) is not unique up to a constant. Indeed, in the case $d=m=1$, any function $f:=f_{\varphi}+f_{b,\sigma}$ with $f_{b,\sigma}^{\prime}(x)=c_{0}\exp(\int_{0}^{x}-\frac{2b(y)}{\sigma(y)^{2}}\mathrm{d}y)$ for any constant $c_{0}$ is also a solution of Eq. (2.6).
In the rest of this paper, we always take $f_{\varphi}$ as in (2.7).
@@ 2.3. Ergodic Estimate of One-step Approximations
Let $\tau\in(0,1)$ be a fixed step-size, $t_{k}:=\tau k$, and $\delta_{k}W:=W(t_{k+1})-W(t_{k})$, $k\in\mathbb{N}$. We consider the following general one-step approximation of Eq. (1.1):
@@
where the explicit or implicit function $B_{b,\sigma}$ depends on the coefficients $b$ and $\sigma$, the step-size $\tau$, the previous value $Y_{k}$, the current value $Y_{k+1}$, and the Brownian increment $\delta_{k}W$ in the subinterval $t_{k},t_{k+1}]$, $k\in\mathbb{N}$.
Denote by $Y_{k}^{x}$ the $k$-th step iterative starting from $Y_{0}=x$. We assume the discrete process $(Y_{k})_{k\in\mathbb{N}}$ is a homogeneous Markov chain with a unique invariant measure $\pi_{\tau}$.
Assume $(Y_{k})_{k\in\mathbb{N}}$ defined in (2.9) has a continuous interpolation $(\hat{Y}_{t})_{t\geq 0}$ which is an Itô process and has the following form in the subinterval $0,\tau]$:
@@
with $\mathscr{F}_{0}$-measurable r.v.s $\hat{Y}_{0}$ and $\tilde{Y}_{0}$ depending on the data of Eq. (1.1). More precisely, we assume that there exists two functions $g_{\tau},\tilde{g}_{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\tilde{Y}_{0}=\tilde{g}_{\tau}(Y_{0})$.
Denote by $\hat{Y}_{\tau}^{x}$ the solution $\hat{Y}_{\tau}$ of (2.10) with initial datum $\hat{Y}_{0}=x$ and define a linear operator
@@
with $x,y\in\mathbb{R}^{d}$.
We now state the main result on the ergodic error between the exact invariant measure $\pi$ of Eq. (1.1) and the numerical invariant measure $\pi_{\tau}$ of the associated one-step approximation (2.9).
@@
Let $X_{0}\sim\pi_{\tau}$ and $\varphi\in\mathcal{B}_{b}(\mathbb{R}^{d})$ or ${\rm Lip}(\mathbb{R}^{d})$ such that $P_{\cdot}\varphi\in L^{1}(0,\infty;\mathcal{C}_{b}^{2}(\mathbb{R}^{d}))$. It follows from the Stein equation (2.6) and the representation (2.12), we have
@@
which proves (2.14). ∎
@@ 3. Applications to Superlinear SODEs
This section will apply the results in Theorem 2.3 to a family of one-step approximations for superlinear SODEs.
Though one can utilize the same methodology to handle a large class of examples, we focus on several recently studied one-step numerical schemes–including the tamed Euler, projected Euler, and backward Euler methods in Example 3.3.
@@ 3.2. Estimates of Stein Equation
We have the following estimates on the directional derivative $\nabla^{i}X_{t}^{\cdot}$ with $i=1,2,3,4$ and $t\geq 0$; see [12, Lemma 3.3] for a weak result in the case of weak dissipativity.
@@
We mention that the analysis of $I_{211},I_{212},I_{213}$ is similar and we take $I_{211}$ as an example. Bt Assumption 3.2, Hölder inequality, and the estimates (3.6) and (3.7), we have
@@
Thus, we conclude (3.3) by the estimates (3.6), (3.7), (3.8), and (3.2).

### v2

@@ Abstract.
This paper establishes the first-order convergence rate for the ergodic errors and ergodic limits of numerical approximations to a class of stochastic ODEs (SODEs) with superlinear coefficients and multiplicative noise. By leveraging the generator approach within the Stein method, we derive a general error representation formula for one-step numerical schemes. Our framework applies to several recently studied schemes, including the tamed Euler, projected Euler, and backward Euler methods. Numerical experiments confirm the optimality of the derived ergodic estimates.
@@
Ergodic estimate, Stein method, stochastic one-step approximations, superlinear stochastic differential equations
@@ 1. Introduction
Here $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are measurable superlinear functions satisfying certain monotone and polynomial growth conditions.
@@
Under appropriate conditions, Eq. (1.1) admits a unique invariant measure $\pi$. Explicitly characterizing $\pi$ is generally infeasible and is known only for specific cases, such as the gradient Langevin equation with additive noise or one-dimensional SODEs with multiplicative nondegenerate noise [6]. Consequently, constructing accurate numerical approximations of $\pi$ is a problem of significant interest. The study of long-time approximations for SODEs is of growing importance, with substantial applications in scientific computing, statistics, machine learning, and generative AI. For instance, sampling in score-based diffusion models – a key technique in generative AI – fundamentally relies on long-time SODE approximations.
@@
The Euler–Maruyama (EM) scheme is effective for Lipschitz coefficients (see, e.g., [2, 17]) or Markovian switching diffusion (see, e.g., [18]). However, for coefficients with superlinear growth, the EM scheme would blow up in the $p$-th moment for all $p\geq 2$ [3], which precludes ergodicity. As a result, several modified schemes have been developed to preserve ergodicity, including the backward Euler method (BEM) [7, 14, 20], the stochastic theta method [10], and modified Euler methods (MEMs) such as the tamed Euler method (TEM) and projected Euler method (PEM) [12, 13]; see also [5] for Markovian switching diffusions.
@@
A central question is to quantify the ergodic error $|\pi(\varphi)-\pi_{\tau}(\varphi)|$ between the numerical invariant measure $\pi_{\tau}$ and the exact invariant measure $\pi$ for a given step-size $\tau\in(0,1)$ and a test function $\varphi$. Using the generator approach [1] within the Stein method [19], the authors of [2] derived an ergodic error representation for additive noise cases. They obtained a convergence rate of $\mathscr{O}(\tau^{1/2}|\ln\tau|)$ for the EM scheme (the unadjusted Langevin algorithm in the Langevin equation case) with Lipschitz drift and nondegenerate additive noise, where $\varphi$ is a Lipschitz function. This result was later generalized in [4] to dissipative Lipschitz systems with nondegenerate multiplicative noise and jump diffusions, removing the logarithmic factor for test functions in $\mathcal{C}_{b}^{3}(\mathbb{R}^{d})$. We also refer to [16] and the references therein for a nonasymptotic bound analysis of the decreasing-step EM scheme.
In practical applications, a more natural goal is to approximate the ergodic limit, i.e., the spatial average $\pi(\phi)$ for a given test function $\phi$, using time-averaging estimators of the form $\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k})$. Such estimators can be effectively computed via simulation. One then aims to characterize the error $|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k})-\pi(\phi)|$ with respect to the number of steps $N$ and the step-size $\tau$. The authors of [15] derived error estimates for time-averaging estimators of numerical schemes, based on the associated Stein equation and an assumption of local weak convergence order, for Eq. (1.1) with Lipschitz coefficients on a torus (so that all coefficients are uniformly bounded). The superlinear case remains unknown. Our representation of the ergodic error $|\pi(\varphi)-\pi_{\tau}(\varphi)|$ leads to a sharp error estimate for the ergodic limit, provided that geometrical ergodicity holds for the general class of one-step numerical approximations considered.
@@
Since the weak convergence rate is typically twice the strong convergence rate, an ergodic error of order $\mathscr{O}(\tau)$ and an ergodic limit error of order $\mathscr{O}(\tau+N^{-1})$ are expected. This paper achieves this optimal first-order rate for a family of superlinear SODEs with multiplicative noise and a class of one-step numerical approximations, assuming test functions belong to $\mathcal{C}_{b}^{4}(\mathbb{R}^{d})$.
To derive the above optimal convergence rates for the ergodic error and ergodic limit for a general one-step approximation, we first derive the Stein equation associated with the invariant measure of Eq. (1.1), as in [2], and present one solution of this equation in terms of the Markov semigroup and the invariant measure (see Theorem 2.1).
By exploiting the stationary property of invariant measures, we establish a relation between the discrete generator of the numerical approximations and the generator of Eq. (1.1). This allows us to derive a new representation of the ergodic error for a general class of one-step approximations to Eq. (1.1) (see Theorem 2.3). This error representation consists of two parts: the first pertains to the error induced by the choice of initial datum, and the second concerns the remainder error arising from the continuous interpolation process associated with the one-step approximations. An error estimate for the ergodic limit follows immediately, provided that geometrical ergodicity holds for these approximations.
Finally, we demonstrate that a general one-step approximation naturally fits into the aforementioned framework. In particular, we establish the optimal first-order convergence of the ergodic error for a sequence of numerical approximations, including the recently investigated TEM, PEM, and BEM, applied to the superlinear SODE (1.1) driven by multiplicative noise. This analysis requires boundedness of the fourth derivatives for the solution of the Stein equation, which is ensured by the integrability in $\mathbb{R}_{+}$ of the fourth derivatives for the solution of Eq. (1.1) with respect to the initial datum.
@@
The paper is organized as follows. Section 2 presents the main ideas, including the Stein equation and the general ergodic error representations. Section 3 applies these results to superlinear SODEs and provides detailed analyses for specific numerical schemes. Numerical experiments verifying the optimality of these ergodic estimates are presented in the final section.
@@ 2. Main Idea and Results
This section introduces frequently used notations, the Stein equation and its solution, and the main result on the ergodic error for general one-step approximations.
@@ 2.1. Notations
Let us begin with some frequently used notations. Throughout, $\left\lvert\cdot\right\rvert$, $\langle\cdot,\cdot\rangle$, and $\left\lVert\cdot\right\rVert_{HS}$ denote the Euclidean norm, the inner product in $\mathbb{R}^{d}$, and the Hilbert–Schmidt norm in $\mathbb{R}^{d\times m}$, respectively.
We use $\mathcal{B}_{b}(\mathbb{R}^{d})$ and ${\rm Lip}(\mathbb{R}^{d})$ to denote the Banach spaces of all bounded measurable and Lipschitz continuous mappings $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, for $k\in\mathbb{N}_{+}$, let $C_{b}^{k}(\mathbb{R}^{d})$ be the Banach space of all $k$-times continuously differentiable functions with bounded partial derivatives $\nabla^{i}\varphi(x)\in\mathcal{L}((\mathbb{R}^{d})^{\otimes i};\mathbb{R})$, the $i$-fold linear functional in $\mathbb{R}^{d}$ with $1\leq i\leq k$, endowed with the seminorm
@@ 2.2. Stein Equation
Assume that $(X_{t}^{\cdot})_{t\geq 0}$ has bounded algebraic moments in the sense that
@@
and that $\mathbb{E}\int_{0}^{T}|\sigma_{j}(X_{t}^{\cdot})|^{2}\mathrm{d}t<\infty$ for all $T>0$ and $j=\{1,2,\cdots,m\}$, ensuring that $\int_{0}^{\cdot}\sigma(X_{s}^{\cdot})\mathrm{d}W(s)$ is a martingale. Note that the definition (2.1) is also valid for $\varphi\in{\rm Lip}(\mathbb{R}^{d})$. Assume that $(P_{t})_{t\geq 0}$ has a unique invariant measure $\pi$ with bounded algebraic moments and the mixing property:
@@
For any $f\in C^{2}(\mathbb{R}^{d})$, Itô formula gives
@@
If $X_{0}\sim\pi$, then $\int_{0}^{t}\mathbb{E}[\mathcal{A}f(X_{s})]\mathrm{d}s=0$ for any $t\geq 0$. By the continuity of $X$, this implies $\mathbb{E}[\mathcal{A}f(X_{0})]=0$.
@@
which shows that $f_{\varphi}$ given in (2.7) solves the Stein equation (2.6). ∎
@@
The solution to the Stein equation (2.6) is not unique even up to a constant. Indeed, in the case $d=m=1$, any function $f:=f_{\varphi}+f_{b,\sigma}$ with $f_{b,\sigma}^{\prime}(x)=c_{0}\exp(\int_{0}^{x}-\frac{2b(y)}{\sigma(y)^{2}}\mathrm{d}y)$ for any constant $c_{0}$ is also a solution of Eq. (2.6).
In the rest of this paper, we always take $f_{\varphi}$ as defined in (2.7).
@@ 2.3. Ergodic Estimate of One-step Approximations
Let $\tau\in(0,1)$ be a fixed step-size, $t_{k}:=\tau k$, and $\delta_{k}W:=W(t_{k+1})-W(t_{k})$ for $k\in\mathbb{N}$. We consider the following general one-step approximation of Eq. (1.1):
@@
where the explicit or implicit function $B_{b,\sigma}$ depends on the coefficients $b$ and $\sigma$, the step-size $\tau$, the previous value $Y_{k}$, the current value $Y_{k+1}$, and the Brownian increment $\delta_{k}W$ over the subinterval $t_{k},t_{k+1}]$, $k\in\mathbb{N}$.
Denote by $Y_{k}^{x}$ the $k$-th iterative starting from $Y_{0}=x$. We assume that the discrete process $(Y_{k})_{k\in\mathbb{N}}$ is a homogeneous Markov chain with a unique invariant measure $\pi_{\tau}$.
Assume that $(Y_{k})_{k\in\mathbb{N}}$ defined in (2.9) has a continuous interpolation $(\hat{Y}_{t})_{t\geq 0}$ which is an Itô process and has the following form on the subinterval $0,\tau]$:
@@
with $\mathscr{F}_{0}$-measurable random variables $\hat{Y}_{0}$ and $\tilde{Y}_{0}$ depending on the data of Eq. (1.1). More precisely, we assume there exists two functions $g_{\tau},\tilde{g}_{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\tilde{Y}_{0}=\tilde{g}_{\tau}(Y_{0})$.
Denote by $\hat{Y}_{\tau}^{x}$ the solution $\hat{Y}_{\tau}$ of (2.10) with initial datum $\hat{Y}_{0}=x$, and define a linear operator
@@
for $x,y\in\mathbb{R}^{d}$.
We now state the main result on the error representation between the exact invariant measure $\pi$ of Eq. (1.1) and the numerical invariant measure $\pi_{\tau}$ of the associated one-step approximation (2.9), along with the corresponding error estimate for the ergodic limit.
@@
Assume furthermore that the one-step approximation (2.9) is $V$-geometrically ergodic (polynomially mixing), i.e., there exist positive constants $C$ and $\rho\in(0,1)$ such that $|\mathbb{E}\varphi(Y_{k}^{x})-\pi_{\tau}(\varphi)|\leq CV(x)\rho^{k}$. Then
$\Big|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k}^{x})-\pi(\varphi)\Big|$
(2.15) $\leq\frac{CV(x)}{(1-\rho)N}+\tau^{-1}\Big|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(g_{\tau}(X_{0}))-\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=X_{0}}]\Big|.$
@@
Let $X_{0}\sim\pi_{\tau}$ and $\varphi$ as stated. From the Stein equation (2.6) and the representation (2.12), we have
@@
which proves (2.14).
By triangle inequality and the $V$-geometrical ergodicity of $\{Y_{k}\}$, we obtain
$\Big|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k}^{x})-\pi_{\tau}(\varphi)\Big|$ $\leq\Big|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k}^{x})-\pi_{\tau}(\varphi)\Big|+|\pi_{\tau}(\varphi)-\pi(\varphi)|$
$=\frac{1}{N}\sum_{k=0}^{N-1}|\mathbb{E}\varphi(Y_{k}^{x})-\pi_{\tau}(\varphi)|+|\pi_{\tau}(\varphi)-\pi(\varphi)|.$
Inequality (2.3) then follows from the above estimate and the representation (2.14). ∎
@@ 3. Applications to Superlinear SODEs
This section applies the results of Theorem 2.3 to a family of one-step approximations for superlinear SODEs.
Although the same methodology can handle a broad class of examples, we focus on several recently studied one-step numerical schemes, including the tamed Euler, projected Euler, and backward Euler methods, in Example 3.3.
@@ 3.2. Estimates of Stein Equation
We have the following estimates on the directional derivatives $\nabla^{i}X_{t}^{\cdot}$ with $i=1,2,3,4$ and $t\geq 0$; see [12, Lemma 3.3] for a weak result in the case of weak dissipativity.
@@
We mention that the analysis of $I_{211},I_{212},I_{213}$ is similar and we take $I_{211}$ as an example. By Assumption 3.2, Hölder inequality, and the estimates (3.6) and (3.7), we have
@@
Thus, we obtain the estimate (3.3) from (3.6), (3.7), (3.8), and (3.2).